OpenStudy (anonymous):

Could someone please explain to me how they got this answer? (√2/2)/(1+√2/2) = √2/(2+√2)

OpenStudy (zehanz):

First, let's try to write the left side of the formula in the proper way: \(\dfrac{\dfrac{\sqrt{2}}{2}}{1+\dfrac{\sqrt{2}}{2}}\), or is it: \(\dfrac{\dfrac{\sqrt{2}}{2}}{\dfrac{1+\sqrt{2}}{2}}\)?

OpenStudy (anonymous):

The first one is correct.

OpenStudy (jdoe0001):

\(\bf \cfrac{ \frac{\sqrt{2}}{2} }{ 1+\frac{\sqrt{2}}{2} }\qquad recall \implies \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}\qquad thus \\ \quad \\ \cfrac{ \frac{\sqrt{2}}{2} }{ 1+\frac{\sqrt{2}}{2} }\implies \cfrac{ \frac{\sqrt{2}}{2} }{ \frac{2+\sqrt{2}}{2} }\implies \cfrac{\sqrt{2}}{2}\cdot \cfrac{2}{2+\sqrt{2}}\implies \cfrac{\cancel{2}\sqrt{2}}{\cancel{2}[2+\sqrt{2}]}\)

OpenStudy (zehanz):

OK, the double fractions are ugly, so we will try to get rid of them. Why not multiply numerator and denominator with 2?

OpenStudy (zehanz):

Wait, I see that my good friend @jdoe0001 has done all the work for me (and you). :( Wouldn't it be better to leave something to do for @September1620?

OpenStudy (anonymous):

I will actually have to write it all out following his work to understand it. If I just wanted the answer to my homework it's already been graded, I'm just trying to understand where I went wrong. Thanks.

OpenStudy (zehanz):

YW!